OpenStudy (anonymous):

Find −a2 − 3b3 + c2 + 2b3 − c2 if a = 3, b = 2, and c = −3.

OpenStudy (anonymous):

the answer is -12

OpenStudy (anonymous):

How did you got -12

OpenStudy (anonymous):

i plugged in the numbers. ill go ahead and type the equation after the multiplication..

OpenStudy (anonymous):

-6-18+(-6)+12-(-6)

OpenStudy (anonymous):

-6-18 = -24 -24+(-6) = -30 -30 +12 = -18 -18 -(-6) --> -18+6 = -12